What speed would the motorist have to average on the return trip from B to A to yield an overall speed of 30km per hour ( assume the motorist has average 20km per hour from A to B)
1
Expert's answer
2020-06-10T19:54:06-0400

We denote by "l" the distance between A and B. Then for the trip from A to B he spent "\frac{l}{20}" hours. We denote by "v" his average speed on the return trip. The time he used on the return trip is "\frac{l}{v}" hours. Since the overall speed of the motorist is 30 km per hour, we get the equation for "v":

"30=\frac{2l}{\frac{l}{20}+\frac{l}{v}}\,\,\Longrightarrow\,\,30=\frac{40v}{20+v}\,\,\Longrightarrow
600=10\,v\,\,\Longrightarrow v=60 \,\,\textrm{km\,\, per\,\, hour}"

Answer: 60 km per hour